OpenStudy (anonymous):

If f(x) = 2x2 + 4, find f(3). 22 16 −0.5 40

OpenStudy (anonymous):

I got 12

OpenStudy (anonymous):

I wrote that out wrong.

OpenStudy (anonymous):

2X^2+4

OpenStudy (anonymous):

f(3) = 2(3)^2 +4 f(3) = 2(9) + 4 f(3) = 18+4 f(3) = 22